OpenStudy (anonymous):

Can anyone help with sequences? Medal to who can help. Find an explicit rule for the nth term of the sequence. 2, -8, 32, -128, ... answer choices: an = 2 ⋅ 4^n+1 an = 2 ⋅ (-4)^n an = 2 ⋅ 4^n-1 an = 2 ⋅ (-4)^n-1 n-1 means number before what ever n is right? n means the number? I am confused. Thanks to anyone who can help

OpenStudy (amistre64):

think of n as "now"; and n-1 as "then" so yes

OpenStudy (anonymous):

well N means the Nth term. Yes n-1 means the number before the number n. The answer here would be D. Example - A1 = 2. Now for A2 if we use An = 2.(-4)^n-1, we have A2 = 2.(-4)^2-1 which is 2.(-4)^1 = 2.-4 = -8. Similarly A3 = 2.(-4)^3-1 => 2.(-4)^2 = 2.16 = 32.

OpenStudy (anonymous):

OH! so, N is the position of the number in the sequence?

OpenStudy (anonymous):

Because if 2 was position 1 1-1 would be 0 2-0 would be 2 which is the first term and so on!

OpenStudy (amistre64):

if the sequence starts with n=1, then yes

OpenStudy (anonymous):

Thanks so much guys!